OpenStudy (anonymous):

If f(x) = 2 x^{2 x}, find f'( 2 ).

OpenStudy (anonymous):

\[If f(x) = 2 x^{2 x}, find f'( 2 )\]

OpenStudy (pradius):

dont u need to only plug the 2 into the x?

OpenStudy (anonymous):

Yah after I find the derivativve wwhich I dont know how

Parth (parthkohli):

No, you have to find the derivative and then plug 2 into that.

OpenStudy (aripotta):

what's a derivative?

OpenStudy (anonymous):

Can someone help me find the derivative

Parth (parthkohli):

Okay...\[{d \over dx} 2x^{2x} \implies 2{d \over dx}x^{2x}\]Chain rule now

OpenStudy (anonymous):

Okay so is it 2x*2x*2 ?

OpenStudy (anonymous):

4*4*2?

Parth (parthkohli):

Erm, no.

OpenStudy (anonymous):

dang it!

Parth (parthkohli):

You have to use the Chain Rule, Monique.

OpenStudy (anonymous):

I thThere is no exponent rule?

OpenStudy (anonymous):

ohh I pull the 2 out

Parth (parthkohli):

There's no power rule for this one, but you can pull the constant coefficients out, yes.

hartnn (hartnn):

this is an example of log diff. let I= x^{2x} take log on both sides, what u get ?

OpenStudy (anonymous):

4x*x*2?

OpenStudy (anonymous):

oh wait

OpenStudy (anonymous):

darn it

Parth (parthkohli):

You can use @hartnn's method too. I learnt it that way, originally!

Parth (parthkohli):

I think that you are comfortable with @hartnn's method if you understand what logarithms are.

OpenStudy (anonymous):

lny=2xlnx

OpenStudy (anonymous):

Either use @hartnn method by solving for g(x) = ln(f(x)) and finding first g`(x), or use chain rule by writing f(x,y)=x^y and y=2x, than diferentiate.

OpenStudy (anonymous):

Still lost

Parth (parthkohli):

Let's turn out with something easy ;)\[a^b = c \implies \log_a c =b\]\[x^{2x} = l\implies \log_x l = 2x\]

hartnn (hartnn):

CANNOT USE CHAIN RULE HERE only log diff.

OpenStudy (anonymous):

Help me!!!!

hartnn (hartnn):

I = x^(2x) log I = 2x log x .....(log here is natural log ln) ok with 1st step ?

OpenStudy (anonymous):

what is the l?

OpenStudy (anonymous):

lny=2x ln x

hartnn (hartnn):

I is anything... i mean u can take y also so yes, ln y = 2x ln x now u know if u now differentiate, how would u diff ln y ??

OpenStudy (anonymous):

Im a little confuesed on how to differentiate this

hartnn (hartnn):

ok, \(\huge \frac{d}{dx}lny=\frac{1}{y}\quad\frac{dy}{dx}\) ok ?

hartnn (hartnn):

can i assume u can differentiate the right side using product rule ??